Abey​ Kuruvilla, of Parkside​ Plumbing, uses 1,190 of a certain spare part that costs $27 for each​ order, with an annual holding cost of $26.

​a) Calculate the total cost for order sizes of​ 25, 40,​ 50, 60, and 100 ​(round your responses to two decimal​ places).

order size total cost
25 ?
40 ?
50 ?
60 ?
100 ?

b) What is the economic order​ quantity?_____ units ​(round your response to two decimal​ places).

Please
Answer #1

Annual demand(D) = 1190 units

Ordering cost(S) = $27

Holding cost(H) = $26

a) order size = Q

For Q=25,

  • Annual ordering cost = (D/Q) S = (1190/25) 27 = $1285.2
  • Annual holding cost = (Q/2) H = (25/2) 26 = $325.0
  • Total cost = $1285.2+$325.0 = $1610.20

For Q = 40

  • Annual ordering cost = (D/Q) S = (1190/40) 27 = $803.25
  • Annual holding cost = (Q/2) H = (40/2) 26 = $520
  • Total cost = $803.25 + $520 = $1323.25

For Q = 50

  • Annual ordering cost = (D/Q) S = (1190/50) 27 = $642.6
  • Annual holding cost = (Q/2) H = (50/2) 26 = $650
  • Total cost = $642.6 + $650 = $1292.60

For Q = 60

  • Annual ordering cost = (D/Q) S = (1190/60) 27 = $535.5
  • Annual holding cost = (Q/2) H = (60/2) 26 = $780
  • Total cost = $535.5 + $780 = $1315.50

For Q = 100

  • Annual ordering cost = (D/Q) S = (1190/100) 27 = $321.3
  • Annual holding cost = (Q/2) H = (100/2) 26 = $1300
  • Total cost = $321.3 + $1300 = $1621.30

b) Economic order quantity = √(2DS/H)

= √[(2 X 1190 X 27) / 26]

= √(64260/26)

= √2471.54

= 49.71 units